OpenStudy (anonymous):

if I assume that cos(x)+5sin(x)=23/5 and 2cos(x)+sin(x)=10/5 how would I find the exact numeric value of sec(x) and tan(x)?

OpenStudy (asnaseer):

think of cos(x) and sin(x) as two unknowns - call them 'c' and 's' for example, so we have: c = cos(x) s = sin(x) then write the equations using these new symbols as follows: c + 5s = 23/5 2c + s = 10/5 you then have two equations with two unknowns which you should be able to solve to find c and s. then you can use these to get the required values as follows: sec(x) = 1/cos(x) = 1/c tan(x) = sin(x)/cos(x) = s/c